# Accelerated Time-domain analysis for Gravitational Wave Astronomy

### 06 Mar 2026 v1

@@ I Introduction
The plan of the paper is as follows. First, in Sec. II, we describe the statistical basis of the time-domain approach and the whitened representation. We also describe some of the salient features of the pipeline, including handling gaps and finite-segments, which is seamless in the time-domain representation. In Sec. III, we carry out an algorithmic complexity analysis and discuss the feasibility of current and future gravitational wave signals in the context of modern computational hardware. In Sec. IV we describe potential speedups, both algorithmic and hardware, that make time-domain methods feasible for gravitational wave astronomy. In Sec. V, we demonstrate the time-domain pipeline using injections. We also validate and demonstrate its capabilities by applying it to the loudest GW event observed so far, GW250114. We summarize the results in Sec. VI.
@@ II Data model and time-domain likelihood
where $\bm{h}$ is the detector response to a CBC waveform parameterized by $\bm{\theta}$, and $\bm{n}$ the noise. Both can be represented in the form of row vectors of size $N$. Each time-sample of these vectors will be denoted by a Latin index, e.g., $n_{j}$ for the jthjth noise sample, with $j\leq N$.
@@
The Spectral theorem in linear algebra  [35, 72, 19] dictates that a real symmetric matrix admits an eigen-basis in which it is diagonal. If one further imposes periodic boundary conditions on the domain (and thus on the noise and signal), it can be shown that this basis is the Fourier Basis, and that $\bm{C}_{ij}$ becomes circulant, leading to the familiar FD approach, where efficient FFT algorithms are available. In reality, the data and the random process is not periodic. One circumvents this by applying window functions, which taper the data to zero smoothly at both ends. This enables one to make the circulant assumption and reduce spectral leakage. The latter is important in making the inner product in Sec. II stable against small perturbations in the domain. However, the downside is that it attenuates the signal across time samples, as windows have non-local support in the time domain, making it difficult to analyze localized time segments with sharp cutoffs without special techniques [68] or by compromising information at the boundaries.
In what follows, we relax the circulant assumption and remain in the time domain, thereby removing the need for tapering or window functions and enabling the analysis of sharply terminated segments of the signal that start and/or end abruptly.
@@ II.2 Implementation details
Figure 3: The estimated autocorrelation and PSD around the GW230814 event.
@@
    Multi-detector likelihood. The time-domain likelihood model for each detector is initialized as an instance of the ‘GWTransientTD‘ detector class, which encapsulates the data, the whitening operator, methods for waveform generation, and an interface to stochastic samplers such as bilby and dynesty. Multiple detectors are handled by a child class of the ‘GWTransientTD‘ class, which is also composed with $n_{d}$ instances of the ‘GWTransientTD’ class, one for each detector. Each call to the multi-detector instance’s likelihood method iteratively calls the corresponding functions of the individual detectors, and the log-likelihoods are summed. There is also a consistency test infrastructure to assess signal consistency across different time segments of the waveforms. We point the reader to  [111] for details. The basic class structure of tdanalysis is summarized in Fig. 2.
@@
This implementation and choice of segmentation enables the analysis of any portion of the time domain signal, be it inspiral, ringdown, merger, or the whole signal, possibly excluding a portion in between, using the corresponding portion of a full IMR waveform. Once the analysis segments are defined by their GPS times, the peak time of the signal determines the portion of the template waveform that lies in the analysis window. We note that this implementation is an improvement over existing implementations for analyzing individual segments of the signal in many ways. First, in existing time domain studies like  [99, 98, 97], some of the extrinsic parameters like sky locations, polarization, and peak time are fixed due to limitations of speed. The accelerations presented in this work enable sampling over all parameters, including the peak time, with judicious resources and in feasible time. In fact, as described in  [111], we show that one can simultaneously analyze multiple segments of data, corresponding to a 25 dimensional parameter space, owing to the accelerations implemented here. As a corollary, the analysis of the ringdown segment can thus include non-linearities as well, with no ambiguities in the choice of start time, as it is also sampled over.
Related models, that analyze the data using portions of the IMR waveforms have been used in the past for ringdown studies. In [25], the post-merger part of the signal from a non-spinning, quasi-circular source is modeled through the corresponding portion of an effective one body waveform, inferring the ringdown frequencies even though the full IMR signal is analyzed, fixing sky locations in some studies. In  [107, 109, 110], the infalling radiation at the outer common dynamical horizon was fit to the post-peak portion of aligned-spin full IMR waveform to infer the parameters of the source, keeping the start time fixed at the peak of the $\ell=2,m=2$ mode. Later  [30, 118, 59] also use the post-peak portion of a full IMR waveform to analyze the post-merger signal in a similar fashion. The sky location and polarization angles are fixed in  [30]. In  [118], the start time and extrinsic parameters are sampled over, while in  [59], the sky location and peak time are fixed. More recently,  [37] revisit  [30] using modern waveform models with multiple choices of start time, while keeping it fixed for each run. They fix the sky location and the merger time in each of their analysis. Although not addressing consistency across segments,  [99, 98] analyze individual segments of the waveform including the ringdown portion to extract precession physics. They fix the sky position, polarization and the coalescence time in their analysis. However, this implementation allows all parameters to be feasibly sampled over, including the peak time of the signal feasibly.
Secondly, like  [97], this a fully time domain pipeline with a Gaussian Likelihood, thus not needing the Whittle approximation.
@@ III.1.3 Gohlberg - Semencul theorem and Convolution / FFT-based method:
$\bm{L}_{x}$ is the lower-triangular Toeplitz matrix with first column $\mathbf{x}=(x_{1},\dots,x_{n})^{T}$, and $\bm{R}_{x}$ is the strictly upper-triangular Toeplitz matrix with first row $(0,x_{n},x_{n-1},\dots,x_{2})$. This means that evaluating the inner products in (9), which corresponds to evaluating quadratic forms, involves the action of Toeplitz matrices on vectors, which can be circulantly embedded by doubling the lengths of the data vectors. For a Toeplitz matrix $\bm{T}\in\mathbb{R}^{n\times n}$ and vector $\mathbf{x}\in\mathbb{R}^{n}$, let $\mathbf{S}\in\mathbb{R}^{m\times m}$ be a circulant embedding of $\bm{T}$ with $m\geq 2n-1$. If $c$ is the first column of $\mathbf{S}$, and $\tilde{\mathbf{x}}\in\mathbb{R}^{m}$ is x zero-padded to length $m$, then
@@ III.1.4 Reduced-basis compressions:
To summarize this section, various speedup methods improve overall scaling in computational complexity/ memory traffic, yielding a factor of 2 improvement at best, but do not change the overall complexity, which remains at $O(N^{2})$ and $O(MN^{2})$. Only the convolutional FFT-based implementations offer superior scaling similar to native FD methods, at the cost of doubling the effective sample size. In this next subsection, we explore how effective use of hardware could help speedup these algorithms.
@@ III.2 Hardware acceleration
The adaptability of time domain algorithms that are $O(N^{2})$ computationally expensive, as opposed to FFT based methods that are $O(N\log N)$ is primarily due to advancements in computational hardware. This involves enhancements to modern CPU, GPU, and memory over the past two decades. We describe here the specifics of how these improvements enable time-domain methods to be adapted for gravitational-wave astronomy.
@@ III.2.3 GPU acceleration
GPU memory, on the other hand, has also been slowly rising in capacity, reaching up upto 100GB on data centre GPUs and a few GBs on personal machines. This allows matrices of the order 12GBs1-2GBs to be stored on VRAMs, which was infeasible to handle a decade ago on any computing platform. The bandwidth dictates that streaming such matrices takes less than $2ms$, suggesting the same bottleneck, even with GPUs’ improved memory bandwidth. However, on GPUs, there exists another overhead: GPU kernel startup cost. Kernel latency is usually constant, independent of the operation being performed, and is of the order of a few tens of $ms$. Thus, given the requirement for repeated evaluations of the likelihood, and the fact that the template vectors need to be transferred from the CPU at each evaluation, the Kernel latency dictates the speed of evaluations on GPUs, given that the whitening operator $\bm{W}$ and the whitened data vector are cached on VRAM.
@@ V.1 Performance and scaling
Figure 5: Benchmark results for direct Cholesky decomposed whitening operator method, and the GSCE method. The left plot shows the combined number of likelihood calls per second of the sampler across the multiprocessing pool on the left axis, and the single likelihood evaluation time on the right axis. The plot in the centre shows the performance of the GSCE method. The right panel shows the ratio of likelihood evaluation time on the device specified on the x-axis to that of a frequency domain likelihood evaluation on a CPU
We compare the performance of two methods of evaluating the likelihoods i). Using the ‘CDO’ method on CPUs and GPUs, and iii) Using ‘GSCE’ on GPUs. All benchmarks are run on an Intel 258v processor with a memory bandwidth of 8533 MT/s. The device remained plugged into the power socket. The benchmarks are carried out by running the bilby sampler with different numbers of multiprocessing pools, from 1 to 8 to examine the interplay between process based parallelism and shared memory parallelism. At each multiprocessing pool value, the OMP_NUM_THREADS value is set such that OMP_NUM_THREADS $\times npool=8$, the number of physical cores on the device. The results are shown in Fig. 5.
@@
    On GPUs, the ‘CDO’ method attains about 2-10x speedup per likelihood evaluation over CPUs. However, they are still 1-5x slower than the FD method. Using data centre GPUs capable of FP64 compute and optimizing custom GPU kernels with HIP can match CPU FD valuation times.
@@
    Implementing and benchmarking this method, we find that the GSCE method is approximately only slower than FD method by a factor of $\sim 2$. Thus, it is possible to do TD analysis with only twice the computing resources as FD methods.
@@
    On CPUs, the ‘GSCE’ method is roughly 10 times faster than the ‘CDO’ method, while at the same time not using as much memory bandwidth as the ‘CDO’ method, leaving room for process based parallelization. The effective speedup is thus $\approx 10\times n_{pool}$, where $n_{pool}$ is the pool size for proposals.
@@ V.2.3 Real data
We analyze the real data with detector noise for GW250114. The recovered marginal distributions are shown in the third row of Fig. 8. The recovered maximum likelihood waveforms and the residual distribution is shown in Fig. 7. The residual noise distribution had a Kolmogorov–Smirnov (KS) statistic value of $7.81\times 10^{-3}$ and a p-value of $0.505$ for the probability of getting a residual noise distribution at least as extreme as this from a standard normal. The Anderson–Darling test statistic, which weights the tails more heavily than the KS test, returned a value of 0.328. All these values are consistent with a standard normal distribution of residues. The estimated standard deviation of the whitened residues was $0.988$, indicating that the estimated autocorrelation function overwhitened the data by $1.2\%$.

### 10 Apr 2026 v2

@@ I Introduction
The plan of the paper is as follows. First, in Sec. II, we describe the statistical basis of the time-domain approach and the whitened representation. We also describe some of the salient features of the pipeline, including handling gaps and finite segments, which are seamless in the time-domain representation. In Sec. III, we carry out an algorithmic complexity analysis and discuss the feasibility of current and future gravitational wave signals in the context of modern computational hardware. In Sec. IV we describe potential speedups, both algorithmic and hardware, that make time-domain methods feasible for gravitational wave astronomy. In Sec. V, we demonstrate the time-domain pipeline using injections. We also validate and demonstrate its capabilities by applying it to the loudest GW event observed so far, GW250114. We summarize the results in Sec. VI.
@@ II Data model and time-domain likelihood
where $\bm{h}$ is the detector response to a CBC waveform parameterized by $\bm{\theta}$, and $\bm{n}$ the noise. Both can be represented in the form of row vectors of size $N$. Each time sample of these vectors will be denoted by a Latin index, e.g., $n_{j}$ for the jthjth noise sample, with $j\leq N$.
@@
The Spectral theorem in linear algebra  [35, 72, 19] dictates that a real symmetric matrix admits an eigen-basis in which it is diagonal. If one further imposes periodic boundary conditions on the domain (and thus on the noise and signal), it can be shown that this basis is the Fourier Basis, and that $\bm{C}_{ij}$ becomes circulant, leading to the familiar FD approach, where efficient FFT algorithms are available. In reality, the data and the random process are not periodic. One circumvents this by applying window functions, which taper the data to zero smoothly at both ends. This enables one to make the circulant assumption and reduce spectral leakage. The latter is important in making the inner product in Sec. II stable against small perturbations in the domain. However, the downside is that it attenuates the signal across time samples, as windows have non-local support in the time domain, making it difficult to analyze localized time segments with sharp cutoffs without special techniques [68] or by compromising information at the boundaries.
In what follows, we relax the circulant assumption and remain in the time domain, thereby removing the need for tapering or windowing and enabling analysis of sharply terminated segments of the signal that start and/or end abruptly.
@@ II.2 Implementation details
Figure 3: The estimated autocorrelation and PSD around the GW250114 event in the Hanford detector. The left panel shows the autocorrelation function, estimated by first computing the PSD using a media-averaged Welch method, which is shown in blue in the right panel. The band-limited BayesWave estimated PSD used in the FD analysis is shown in orange.)
@@
    Multi-detector likelihood. The time-domain likelihood model for each detector is initialized as an instance of the ‘GWTransientTD‘ detector class, which encapsulates the data, the whitening operator, methods for waveform generation, and an interface to stochastic samplers such as bilby and dynesty. Multiple detectors are handled by a child class of the ‘GWTransientTD‘ class, which is also composed of $n_{d}$ instances of the ‘GWTransientTD’ class, one for each detector. Each call to the multi-detector instance’s likelihood method iteratively calls the corresponding functions of the individual detectors, and the log-likelihoods are summed. There is also a consistency test infrastructure to assess signal consistency across different time segments of the waveforms. We point the reader to  [111] for details. The basic class structure of tdanalysis is summarized in Fig. 2.
@@
This implementation and choice of segmentation enables the analysis of any portion of the time domain signal, be it inspiral, ringdown, merger, or the whole signal, possibly excluding a portion in between, using the corresponding portion of a full IMR waveform. Once the analysis segments are defined by their GPS times, the peak time of the signal determines the portion of the template waveform that lies in the analysis window. We note that this implementation is an improvement over existing implementations for analyzing individual segments of the signal in many ways. First, in existing time domain studies like  [99, 98, 97], some of the extrinsic parameters like sky locations, polarization, and peak time are fixed due to limitations of speed. The accelerations presented in this work enable sampling over all parameters, including the peak time, with judicious resources and in a feasible time. In fact, as described in  [111], we show that one can simultaneously analyze multiple segments of data, corresponding to a 25-dimensional parameter space, owing to the accelerations implemented here. As a corollary, the analysis of the ringdown segment can thus include non-linearities as well, with no ambiguities in the choice of start time, as it is also sampled over.
Related models that analyze the data using portions of the IMR waveforms have been used in the past for ringdown studies. In [25], the post-merger part of the signal from a non-spinning, quasi-circular source is modeled through the corresponding portion of an effective one-body waveform, inferring the ringdown frequencies even though the full IMR signal is analyzed, fixing sky locations in some studies. In  [107, 109, 110], the infalling radiation at the outer common dynamical horizon was fit to the post-peak portion of aligned-spin full IMR waveform to infer the parameters of the source, keeping the start time fixed at the peak of the $\ell=2,m=2$ mode. Later  [30, 118, 59] also use the post-peak portion of a full IMR waveform to analyze the post-merger signal in a similar fashion. The sky location and polarization angles are fixed in  [30]. In  [118], the start time and extrinsic parameters are sampled over, while in  [59], the sky location and peak time are fixed. More recently,  [37] revisits  [30] using modern waveform models with multiple choices of start time, while keeping it fixed for each run. They fix the sky location and the merger time in each of their analysis. Although not addressing consistency across segments,  [99, 98] analyze individual segments of the waveform, including the ringdown portion, to extract precession physics. They fix the sky position, polarization, and the coalescence time in their analysis. However, this implementation allows all parameters to be feasibly sampled over, including the peak time of the signal feasibly.
Secondly, like  [97], this is a fully time domain pipeline with a Gaussian Likelihood, thus not needing the Whittle approximation.
@@ III.1.3 Gohlberg - Semencul theorem and Convolution / FFT-based method:
$\bm{L}_{x}$ is the lower-triangular Toeplitz matrix with the first column $\mathbf{x}=(x_{1},\dots,x_{n})^{T}$, and $\bm{R}_{x}$ is the strictly upper-triangular Toeplitz matrix with first row $(0,x_{n},x_{n-1},\dots,x_{2})$. This means that evaluating the inner products in (9), which corresponds to evaluating quadratic forms, involves the action of Toeplitz matrices on vectors, which can be circulantly embedded by doubling the lengths of the data vectors. For a Toeplitz matrix $\bm{T}\in\mathbb{R}^{n\times n}$ and vector $\mathbf{x}\in\mathbb{R}^{n}$, let $\mathbf{S}\in\mathbb{R}^{m\times m}$ be a circulant embedding of $\bm{T}$ with $m\geq 2n-1$. If $c$ is the first column of $\mathbf{S}$, and $\tilde{\mathbf{x}}\in\mathbb{R}^{m}$ is x zero-padded to length $m$, then
@@ III.1.4 Reduced-basis compressions:
To summarize this section, various speedup methods improve overall scaling in computational complexity/ memory traffic, yielding a factor of 2 improvement at best, but do not change the overall complexity, which remains at $O(N^{2})$ and $O(MN^{2})$. Only the convolutional FFT-based implementations offer superior scaling similar to native FD methods, at the cost of doubling the effective sample size. In this next subsection, we explore how the effective use of hardware could help speed up these algorithms.
@@ III.2 Hardware acceleration
The adaptability of time domain algorithms that are $O(N^{2})$ computationally expensive, as opposed to FFT-based methods that are $O(N\log N)$, is primarily due to advancements in computational hardware. This involves enhancements to modern CPU, GPU, and memory over the past two decades. We describe here the specifics of how these improvements enable time-domain methods to be adapted for gravitational-wave astronomy.
@@ III.2.3 GPU acceleration
GPU memory, on the other hand, has also been slowly rising in capacity, reaching up upto 100GB on data center GPUs and a few GBs on personal machines. This allows matrices of the order 12GBs1-2GBs to be stored on VRAMs, which was infeasible to handle a decade ago on any computing platform. The bandwidth dictates that streaming such matrices takes less than $2ms$, suggesting the same bottleneck, even with GPUs’ improved memory bandwidth. However, on GPUs, there exists another overhead: GPU kernel startup cost. Kernel latency is usually constant, independent of the operation being performed, and is of the order of a few tens of $ms$. Thus, given the requirement for repeated evaluations of the likelihood, and the fact that the template vectors need to be transferred from the CPU at each evaluation, the Kernel latency dictates the speed of evaluations on GPUs, given that the whitening operator $\bm{W}$ and the whitened data vector are cached on VRAM.
@@ V.1 Performance and scaling
Figure 5: Benchmark results for direct Cholesky decomposed whitening operator method, and the GSCE method. The left plot shows the combined number of likelihood calls per second of the sampler across the multiprocessing pool on the left axis, and the single likelihood evaluation time on the right axis. The plot in the center shows the performance of the GSCE method. The right panel shows the ratio of likelihood evaluation time on the device specified on the x-axis to that of a frequency domain likelihood evaluation on a CPU.
We compare the performance of two methods of evaluating the likelihoods i). Using the ‘CDO’ method on CPUs and GPUs, and iii) Using ‘GSCE’ on GPUs. All benchmarks are run on an Intel 258v processor with a memory bandwidth of 8533 MT/s. The device remained plugged into the power socket. The benchmarks are carried out by running the bilby sampler with different numbers of multiprocessing pools, from 1 to 8, to examine the interplay between process-based parallelism and shared memory parallelism. At each multiprocessing pool value, the OMP_NUM_THREADS value is set such that OMP_NUM_THREADS $\times npool=8$, the number of physical cores on the device. The results are shown in Fig. 5.
@@
    On GPUs, the ‘CDO’ method attains about 2-10x speedup per likelihood evaluation over CPUs. However, they are still 1-5x slower than the FD method. Using data center GPUs capable of FP64 compute and optimizing custom GPU kernels with HIP can match CPU FD valuation times.
@@
    Implementing and benchmarking this method, we find that the GSCE method is approximately only slower than the FD method by a factor of $\sim 2$. Thus, it is possible to do TD analysis with only twice the computing resources as the FD methods.
@@
    On CPUs, the ‘GSCE’ method is roughly 10 times faster than the ‘CDO’ method, while at the same time not using as much memory bandwidth as the ‘CDO’ method, leaving room for process-based parallelization. The effective speedup is thus $\approx 10\times n_{pool}$, where $n_{pool}$ is the pool size for proposals.
@@ V.2 Parameter estimation
Figure 9: The recovered maximum Likelihood waveforms. The left panels show whitened (decorrelated and standardized) data from the Hanford detector, and the right panels from Livingston. The top panel shows the maximum-likelihood waveform recovered from the GW250114 IMR analysis in red, with the faint orange band showing the set of all posterior waveforms. The middle panel shows the same for the ringdown analysis, started 10M after the GPS peak time of the polarizations arriving at the geocentre. The bottom panel shows the corresponding plots from a zero-noise injection study, using the maximum-likelihood waveform from the full IMR analysis.
Figure 10: Matched filter SNR distributions. Here, the posterior distribution on the matched filter SNR obtained using a frequency-domain analysis (bilby) is shown, along with the corresponding distribution from this code. Note that the distributions are centered at their median values, which are mentioned in the legend.
Figure 11: Recovery of parameters from a ringdown injection study. This corner plot shows the one, two, and three sigma highest-density intervals of a 2D-Gaussian distribution. The orange lines show the injected value. The rest of the 6 parameters are included in the analysis, but omitted here for brevity.
@@ V.2.1 Zero noise injection
A ringdown-only analysis was also performed using a GW250114-like signal injected with ‘NRSur7dq4‘. The ringdown portion of the injection was analyzed using the corresponding portion of the same waveform model, successfully and consistently recovering the injection. The analysis began 10M after the peak of the polarizations. The set of all whitened waveforms from the posterior distribution of samples is shown in the bottom panel of Fig. 9.
@@ V.2.3 Real data
We analyze the real data with detector noise for GW250114. We carry out three separate analyses, involving (i) the full IMR signal, (ii) the inspiral-only portion truncated at 200M-200M before merger, and (iii) the ringdown portion $10M$ after the peak of the signal. In Fig. LABEL:fig:. The set of all reconstructed posterior waveforms is shown in the top and middle panels. The corresponding waveforms for the inspiral-only analysis are shown in 7
The recovered marginal distributions are shown in the third row of Fig. 8. Some multi-modality is seen in the chirp mass-distribution, which could be caused by the signal length. The recovered maximum likelihood waveforms and the residual distribution are shown in Fig. 7. The residual noise distribution had a Kolmogorov–Smirnov (KS) statistic value of $7.81\times 10^{-3}$ and a p-value of $0.505$ for the probability of getting a residual noise distribution at least as extreme as this from a standard normal. The Anderson–Darling test statistic, which weights the tails more heavily than the KS test, returned a value of 0.328. All these values are consistent with a standard normal distribution of residues.
A consistent notable feature in the time domain analysis across all the analyses carried out is the recovery of higher SNRs and a higher precision thereof. Fig. 10 shows the 1D histogram of the matched filter SNR from a frequency domain analysis using bilby in blue, and a time domain analysis result in orange. The runs are carried out with the same settings, but with different PSDs. The frequency-domain analysis uses the band-limited LVK-released PSD computed with BayesWave, and the time-domain analysis uses a median-averaged Welch PSD, inverse Fourier-transformed to construct an autocorrelation function. The histograms are centered on the respective median values, as indicated in the legend. The 90% CI for the SNR distributions were found to be $76.187^{+0.075}_{-0.045}$ for FD, and $78.77^{+0.035}_{-0.056}$ for TD analysis, respectively. Their standard deviations were found to be $0.037$ and $0.029$, respectively. Running a Gaussianity test on the time domain whitened residues (right panel of  7) shows that the residues are consistent with a Gaussian, with a The mean of the whitened residues is $0.0021$, and the standard deviation is $\sigma_{\tilde{r}}=0.975$. This shows that the Welch PSD is overestimating the noise. Given that the SNRs scale as $\rho_{true}=\rho_{obs}/\sigma_{\tilde{r}}$, and correcting for this bias, the true suggested median SNR recovered by the time domain code is closer to $80$.
In the ringdown analysis performed using GW250114 with ‘NRSur7dq4‘, the recovered posterior distributions were marginally consistent with those of full IMR analysis. The analysis began $10M$ after the peak of the polarizations. The set of all whitened waveforms from the posterior distribution of samples is shown in Fig. LABEL:fig:rd. Fig. 11 shows a corner plot of select parameters from the full 15D posteriors. The injection values are mostly within the $2\sigma$ contours, with some lying on the edge, or outside of the 90% CI of the corresponding marginal distribution. Given that the analysis uses heavily limited information of the signal past $10M$ after the merger, the recovery seems reasonable.